Starting with 9.00g Na and 8.00 g O2 gas, determine the % yield when Sally got 11.0 g Na2O in the lab?

4 Na + O2 --> 2Na2O


1
Expert's answer
2021-03-03T01:57:41-0500

Molar masses:      

Na - 22.99 g/mol

O2 - 32.00 g/mol

Na2O - 61.98 g/mol

Moles               

Na: 9.00 / 22.99 = 0.39147

O2: 8.00 / 32.00 = 0.25 mol

Na2O: 11.0 / 61.98 = 0.177 mol


Stoichiometry:   

each mole of O2 requires 4 times moles Na: 0.25 x 4 = 1 mol. We have only 0.39147 moles Na (Na - limiting).


Moles of Na2O according to the equation: 0.39147 x 2/4 = 0.1957 mol


Mass of the oxide: 0.1957 x 61.98 = 12.1 g – theoretical yield

11.0 / 12.1 = 0.9091 or 90.91 % - practical yield